OpenStudy (anonymous):

Jonathan wants to add a border around the top of the four walls. He bought 50 feet of border. How many more feet of border does Jonathan need to buy to cover his living room walls? Remember, the room is 12 ½ feet by 15 ¾ feet. 6 feet 6.5 feet 7 feet 8 feet

OpenStudy (anonymous):

Find the perimeter: 25 + 30 + 6/4 = 55 + 1 + 1/2 = 56.5 Subtract 50 from it, and get 6.5 ft.

OpenStudy (anonymous):

thank you